OpenStudy (anonymous):

solve the system using substitution. Put the answer in pair form, (x,y) { x - 2y = -4 { 2x - y = 1

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

consider the first equation x-2y=-4 can u pls tell me what u get when u solve it for x.

OpenStudy (anonymous):

is it 2

OpenStudy (ajprincess):

nope.:( solving for a variable gives that variable on one side and rest of the terms attached to it on the other side. x-2y=-4 if u solve for x then u should have x=somthing As u can -2y is attached to x. to get x alone we need to remove it. To do that add 2y.this will give x-2y+2y=-4+2y x=-4+2y we add 2y to the other side since this is an equation both sides should be equal..getting this>

OpenStudy (anonymous):

ohh okay

OpenStudy (ajprincess):

Nw u have x=-4+2y plug in this value of x in 2x - y = 1 . then u will get 2(-4+2y)-y=1 solve this for y